A population consists of the numbers 2, 4, 8, 10 and 5. Let us list all 

the possible samples of size 3 from this population and construct the sampling 

distribution of the sample mean?


1
Expert's answer
2021-05-04T13:28:32-0400

We have given the population "2,4,8,10" and "5."

Sampling distribution of size 3 can be given as





Sampling distribution of sample mean,

"P(\bar X=4.66) = \dfrac{1}{5} ," "P(\bar X=3.66) = \dfrac{1}{10},"

"P(\bar X=6.66) = \dfrac{1}{10}," "P(\bar X = 7.66) = \dfrac{1}{10},"

"P(\bar X=5.66)= \dfrac{1}{10}"

"P(\bar X=5) = \dfrac{1}{5}"

"P(\bar X=7.33) = \dfrac{1}{10}"

"P(\bar X=6.33) = \dfrac{1}{10}"